Kali and Aliyah are selling pies for a school fundraiser. Customers can buy apple pies and lemon
meringue pies. Kali sold 11 apple pies and 6 lemon meringue pies for a total of $126. Aliyah
sold 12 apple pies and 12 lemon meringue pies for a total of $192. What is the cost each of one
apple pie and one lemon meringue pie?

Answer by Boreal(15235)

11A+6L=126
12A+12L=192
multiply top by -2 and add
-22A-12L=-252
12A+12L=192
-10A=-60
A=$6
1A+1L=16 from equation 2, divided by 12
L=$10